OpenStudy (anonymous):

Solve, using factoring: 8x2 = 16x Answers: {0} {-8, -16} {0, 8} {0, 2}

OpenStudy (anonymous):

Solve for x over the real numbers: 8 x^2 = 16 x Move everything to the left hand side. Subtract 16 x from both sides: 8 x^2-16 x = 0 Factor the left hand side. Factor x and constant terms from the left hand side: 8 x (x-2) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by 8: x (x-2) = 0 Solve each term in the product separately. Split into two equations: x-2 = 0 or x = 0 Look at the first equation: Solve for x. Add 2 to both sides: Answer: | | x = 2 or x = 0

OpenStudy (nincompoop):

nice solution wolframalpha

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

aaa thx, i got the answer already lol

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

8x^(2) = 16x 8x^(2) - 16x = 0 8x(x - 2) = 0 8x = 0 ==> x = 0 or x - 2 = 0 ==> x = 2 x = 0,2